OpenStudy (anonymous):

Find the y-intercept for 3x – 2y = 10 please show work!

OpenStudy (anonymous):

the intercept is when x =0 so if you put x=0 you get -2y=10 then solve for y so it intercepts at y = -5

myininaya (myininaya):

gj kanade

OpenStudy (annon):

3x-10=2y divid by 2 and you get... y=(2/3)x - (10/2) y=mx + b y-int = 10/2 = 5

OpenStudy (annon):

im sorry -5

myininaya (myininaya):

gj annon lol

OpenStudy (annon):

i take to long typing -.-

OpenStudy (anonymous):

-5

OpenStudy (anonymous):

Can someone help me out :) http://openstudy.com/groups/mathematics#/groups/mathematics/updates/4deeea764c0e8b0beee6cfd8